7231/assignments/295420?module item_id- 1084344 (1 point) For each problem, select the best response. (1) Fred keeps his savings in his mattress. He began with $650 trom his mother and adds $100 each year His total savings y after r years are given by the equation OA. y 650 100z OB.y 650+100 Oc. y-100+650z OD. y 100+ O E. None of the above. (2) Smokers dont live as long (on the average) as nonsmokers, and heavy smokers dont live as long as light smokers. You regress the age at death of a group of male smokers on the number of packs per day they smoked. The slope of your regression line O A. will be greater than zero. O B. must be between -1 and 1 OC. will be less than zero. D. cant tell without seeing the data ?? Measi remonte nn vninn rhiliren in Mum hai ndia frim this laset,en aree line fnrnrarietinn hanh iveUserabelhazmiuser balhazmiškeysxM
media%2Fa03%2Fa036b11f-c773-492f-aab6-30
Answer #1

Q1.

Ans: B y = 650 + 100x

The total saving will be the total money accumulated over time. As it was started with 650 Rs and increase 100 after every year. So, after x years, the yearly saving will be 100*x

So total will be 650+100x

Q2.

Option C: Must be less than zero

age_{i,death} = \alpha + \beta X_{i}

That means that coefficient of No. of packs is marginal effect of smoking on life span. The more one will smoke, the less years he will live and hence, the dying age should be low.

So, the coefficient is negative.

So, \beta < 0

Q3.

Ans: Option D: 0.93 cm

Suppose we increase x from x to x + 1

\hat{y_{new}}-\hat{y_{old}} = 6.4 + 0.93(x+1) - (6.4 + 0.93x) = 0.93

So, the marginal effect of arm length on height is 0.93. So unit increase in arm length will affec t height be 0.93

Thanks

Please give me thumbs up if you like, comment if you have any doubts.